Write qn equation of parabola whose directrix is y=3 and vertex (2,2)? Plzz help
Answer by lwsshak3(11628)

Write qn equation of parabola whose directrix is y=3 and vertex (2,2)?
Given parabola opens downwards.
Its basic form of equation: (x-h)^2=-4p(y-k), (h,k)=(x,y) coordinates of the vertex.Given vertex: (2,2)
axis of symmetry: x=2
p=1 (distance from vertex to directrix on the axis of symmetry)
4p=4
Equation of given parabola: (x-2)^2=-4(y-2)